### v2

# Jaike van Twiller1    Yossiri Adulyasak2    Erick Delage2    Djordje Grbic1    Rune Møller Jensen1

Abstract
Reinforcement learning (RL) has shown promise in solving various combinatorial optimization problems. However, conventional RL faces challenges when dealing with real-world constraints, especially when action space feasibility is explicit and dependent on the corresponding state or trajectory. In this work, we focus on using RL in container shipping, often considered the cornerstone of global trade, by dealing with the critical challenge of master stowage planning. The main objective is to maximize cargo revenue and minimize operational costs while navigating demand uncertainty and various complex operational constraints, namely vessel capacity and stability, which must be dynamically updated along the vessel’s voyage. To address this problem, we implement a deep reinforcement learning framework with feasibility projection to solve the master stowage planning problem (MPP) under demand uncertainty. The experimental results show that our architecture efficiently finds adaptive, feasible solutions for this multi-stage stochastic optimization problem, outperforming traditional mixed-integer programming and RL with feasibility regularization. Our AI-driven decision-support policy enables adaptive and feasible planning under uncertainty, optimizing operational efficiency and capacity utilization while contributing to sustainable and resilient global supply chains.
1 Introduction
In recent years, machine learning (ML) for combinatorial optimization (CO) has gained much traction Bengio et al. (2021). Learning advanced planning policies yields promising results in solving CO problems in transportation and logistics. Such approaches can outperform existing solution methods in some well-known problems, such as vehicle routing Kool et al. (2019); Hottung et al. (2022) or job shop scheduling Kwon et al. (2020). Even though it is important to benchmark algorithms on traditional CO problems, a significant opportunity remains to address lesser-known yet critical planning challenges in supply chains. By addressing these challenges, we bridge the gap between theory and practice, enhancing the reliability and efficiency of worldwide supply chain operations.
Container shipping is a key component of the global supply chain, responsible for transporting 45% of annual goods valued at \$8.1 trillion UNCTAD (2021). Due to its scale, it is often regarded as the cornerstone of worldwide trade and modern consumerism. However, this scale also has a significant environmental impact, with yearly CO2 emissions exceeding 200 million tonnes Lloyd’s List (2022). Despite this, container vessels are considered an environmentally friendly mode of transportation, due to their relatively low emissions per cargo ton-mile Jensen et al. (2018). Consequently, container shipping is recognized to play a crucial role in the global green transition European Commission (2023). Within container shipping, there are several interdependent and complex planning tasks, such as berth allocation Martin-Iradi et al. (2022), pre-marshalling Hottung et al. (2020), quay crane scheduling Herup et al. (2022), and container vessel stowage planning Van Twiller et al. (2024b). Due to decision-making under demand uncertainty and various combinatorial aspects - such as capacity limits, seaworthiness requirements, minimizing overstowage and maximizing revenue - stowage planning is particularly challenging. To mitigate this challenge, stowage planning is often decomposed into the master planning problem (MPP), which involves the assignment of cargo to clusters of slots, and the slot planning problem (SPP), which allocates containers to individual slots Pacino et al. (2011). However, these subproblems in representative form, particularly the MPP, remain non-trivial and difficult to solve especially in the presence of uncertainty Van Twiller et al. (2024b).
In addition to its complexity, stowage planning faces public data scarcity and relies heavily on human planners supported by limited decision support systems Jensen et al. (2018). Planners should maximize capacity utilization while managing demand uncertainty by accounting for multiple future scenarios. However, evaluating such scenarios is computationally intractable, due to the problem’s complexity and dynamic nature. These limitations highlight the need for efficient, adaptive, and feasible decision-support systems, offering an opportunity for AI-driven solution methods to enhance the resilience and sustainability of the global supply chain.
This paper presents a deep reinforcement learning (DRL) approach with feasibility projection to construct adaptive and feasible solutions, providing a decision-support policy for the MPP under demand uncertainty.
Our main contributions are as follows:
  • Master Stowage Planning Environment: We develop a novel Markov decision process (MDP) for master stowage planning under demand uncertainty, incorporating realistic problem-specific constraints. To address data scarcity, we release the environment as an open-source implementation [^2] https://github.com/OptimalPursuit/navigating_uncertainty_in_mpp.
  • Feasibility Projection: We incorporate differentiable projection layers, including weighted scaling, policy clipping, and violation projection, to enforce inequality constraint satisfaction in DRL frameworks.
  • Efficient and Adaptive Solutions: Our experiments demonstrate that our policy efficiently generates adaptive and feasible solutions under demand uncertainty, significantly outperforming well-known DRL methods and a multi-stage stochastic MIP model.
  • Decision Support for Stowage Planning: Our decision-support policy transcends deterministic models, enabling dynamic and uncertainty-informed planning in a critical part of the global supply chain.
2 Domain Preliminaries
We present the preliminaries of the stowage planning domain to provide the planning context. We refer to Jensen et al. (2018) for a comprehensive review.
Voyages. Liner shipping companies utilize fleets of container vessels that operate on fixed schedules along a closed-loop course, similar to a maritime bus service that carries goods between ports. The complete journey covering all ports is called a voyage, typically starting in regions with a supply surplus (e.g., Asia) and navigating towards high-demand discharge areas (e.g., Europe), often voyages cover more than ten ports. A voyage can be described as a directed path graph $G_{P}=(P,E_{P})$ with nodes being $P=\{1,2,\dots,N_{P}\}$, and edges being legs between ports $E_{P}=\{(p,p\!+\!1)\mid p\in\{1,2,\dots,N_{P}\!-\!1\}\}$ with $N_{P}$ being the last port. We also define sub-voyages by set ${P}_{\text{start}}^{\text{end}}=\{p\in{P}\!\mid\!\text{start}\leq p\leq\text{end}\}$.
Containers have a port of load (pol) and a port of discharge (pod), also named transports or origin-destination pairs. Consider $\textit{pol}\in{P}_{1}^{N_{P}-1}$ and $\textit{pod}\in{P}_{2}^{N_{P}}$ with $P^{2}=P\times P$, which can be represented by transport $\textit{tr}=(\textit{pol},\textit{pod})$. We can define a set of transports $\textit{TR}=\{(i,j)\in{P}^{2}\!\mid\!i<j\}$.
[…]
Vessel Layout. Figure 1 shows the cellular layout of container vessels with standardized coordinates of (bay, row, tier). It is divided into bays (02-38) ranging from the front (e.g., bay 02) to the rear (e.g., bay 38), also named from fore to aft. Bays are defined by an ordered set ${B}=\{1,2,\dots,N_{B}\}$, ranging from fore to aft with $N_{B}$ being the last bay. Each bay contains stacks of cells, whereas cells contain two slots that can hold two 20-foot containers or one 40-foot container. Bays are horizontally separated by hatch covers into above and below deck sections, introducing the set of decks ${D}=\{d^{\textit{above}},d^{\textit{below}}\}$. It is worth noting that hatch covers introduce complexity since below-deck cargo can only be discharged if the above-deck cargo is cleared.
Figure 1: The side and top view of a container vessel.
Utilization. Vessel utilization $u_{p}\in\mathbb{R}^{|B|\times|D|\times|K|\times|\textit{TR}|}$ represents the cargo placement across bays $B$, decks $D$, cargo types $K$, and transports TR. Load operations are defined by $u^{+}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$, whereas discharging is denoted by $u^{-}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$. The utilization at port $p$ is defined as $u_{p}=u_{p-1}+u^{+}_{p}-u^{-}_{p}\;\forall p\in{P}$, where $u_{0}$ is the vessel’s arrival condition at the first port. We also define the vessel’s pre-loading utilization $u^{\prime}_{p}=u_{p-1}-u^{-}_{p}$.
Objectives and Global Constraints. The MPP aims to create a capacity management plan for each general location (indexed by bay and deck) while adhering to global constraints, as outlined below. The MPP maximizes revenue by optimizing capacity for high-value cargo and minimizes costs through improved operational efficiency. The SPP then assigns containers to slots within the general locations while considering local constraints. Due to frequent re-planning as new information arises, algorithmic runtimes over 10 minutes are impractical for real-world operations Pacino et al. (2011).
Vessels have different types of capacity related to TEUs, height, weight and special cargo. The amount to which this can be utilized depends on the arrival condition of vessels and uncertain cargo demand. It is therefore non-trivial to ensure optimal utilization of vessel capacities. The key efficiency objectives are minimizing overstowage, an NP-hard task Tierney et al. (2014), and minimizing the quay crane makespan. Overstowage occurs when containers stacked on top block access to those below in need of discharging, requiring additional crane operations to remove and reload the obstructing containers, thereby increasing the workload. In addition, the crane makespan considerably impacts the port stay duration, while ports set makespan targets and overruns incur costs. It is thus beneficial to minimize the makespan by maximizing the average use of parallel cranes, thereby minimizing idle time.
Furthermore and most importantly, vessels must adhere to safety requirements to ensure seaworthiness, which is heavily influenced by weight distribution. Before introducing the constraints, we first clarify that $\otimes$ represents the outer product and $\odot$ denotes the element-wise Hadamard product, with broadcasting applied when shapes are compatible. The longitudinal (lcg) and vertical centres of gravity (vcg), which are associated with trim and metacentric height, of the load $u_{p}$ prior to leaving port $p$ must remain within specified bounds ($\underline{\textit{lcg}},\overline{\textit{lcg}}\in\mathbb{R}_{>0}$), as shown in Constraints (1) and (2). To compute the stability, we define the longitudinal moment $\textit{lm}=\textit{ld}\otimes w$ and vertical moment $\textit{vm}=\textit{vd}\otimes w$, where $w\in\mathbb{R}^{|K|\times|\textit{TR}|}_{>0}$ represents container weights, and $\textit{ld},\textit{vd}\in\mathbb{R}^{|B|\times|D|}$ denote the longitudinal and vertical distances of positions from the vessel’s centre of gravity.
$\underline{\textit{lcg}}\leq\frac{\mathbf{1}^{\top}\big{(}\textit{lm}\odot u_{p}\big{)}}{\mathbf{1}^{\top}\big{(}w\odot u_{p}\big{)}}\leq\overline{\textit{lcg}},\qquad\forall p\in P$ (1)
$\underline{\textit{vcg}}\leq\frac{\mathbf{1}^{\top}\big{(}\textit{vm}\odot u_{p}\big{)}}{\mathbf{1}^{\top}\big{(}w\odot u_{p}\big{)}}\leq\overline{\textit{vcg}},\qquad\forall p\in P$ (2)
In conclusion, the scale and complexity of the problem pose significant challenges. As a result, most models focus on subsets of realistic constraints to remain tractable, often requiring human intervention to ensure practical usability.
3 Related Work
This section positions our work in the literature on stowage planning, stochastic programming and ML in optimization.
Stowage Planning. To address the complexity of stowage planning, various solution approaches have been applied to different problem formulations, including exact methods Roberti and Pacino (2018), linearly relaxed MIP Pacino et al. (2011), matheuristics Parreño-Torres et al. (2021), population-based metaheuristics Chang et al. (2022), neighbourhood-based metaheuristics Pacino (2018), hybrid frameworks Bilican et al. (2020). Despite these techniques, a recent survey highlights that scalable solutions for representative stowage and MPP problems remain unsolved Van Twiller et al. (2024b). Research has largely focused on deterministic problems, overlooking real-world needs for profit maximization in the face of demand uncertainty.
Stochastic Programming. Decision-making under uncertainty is traditionally approached through stochastic programming, where uncertainty is explicitly represented as a set of discrete scenarios evolving over multiple stages in time, approximating the underlying probability distribution Birge and Louveaux (2011). This process results in a scenario tree, whose computational complexity grows exponentially as $\mathcal{O}(b^{T})$, where $b$ is the branching factor and $T$ is the number of stages. Hence, multi-stage problems with $T>2$ and a reasonably large $b$ are often intractable. Common techniques to overcome this complication are scenario reduction to reduce problem size Römisch (2009), decomposition techniques like Benders decomposition or Lagrangian relaxation to divide the problem Rahmaniani et al. (2017), progressive hedging to enforce non-anticipativity iteratively Boland et al. (2018), or approximation methods, such as stochastic dual dynamic programming Shapiro (2011) and sample average approximation Chen and Luedtke (2022), to enhance tractability while preserving solution quality.
Learning with Hard Constraints. In learning solution heuristics, a challenge arises to ensure feasibility in complex action spaces. This challenge is exacerbated when dealing with the dependence on state variables to define the feasible region. Several deep learning approaches have dealt with learning constraints by backpropagation through (in)equality completion Donti et al. (2021), or differentiable projection layers, such as mapping interior points to boundaries Li et al. (2023), convex programming layers Agrawal et al. (2019), or problem-specific repair layers Chen et al. (2024). Moreover, safe reinforcement learning has dealt with constraints by, e.g., constrained MDPs with a primal-dual approach Ding et al. (2020), soft barriers function Wang et al. (2023), and safety shields Alshiekh et al. (2018).
4 Markov Decision Processes
We present a formal MDP and a decomposed MDP to solve using DRL approaches. Technical details about both MDPs are provided in Appendix A.
4.1 Formal MDP
Let ${\mathcal{M}}=({S},{X},\mathcal{T},{\mathcal{R}},{P}^{N_{P}-1}_{1},\gamma)$ define an episodic discounted MDP representing the MPP, where ${S}$ is a set of states, ${X}$ is a set of actions, $\mathcal{T}:{S}\times{X}\rightarrow\Delta({S})$ is the transition function, ${\mathcal{R}}:{S}\times{X}\times{P}^{N_{P}-1}_{1}\rightarrow\mathbb{R}$ is the reward function, ${P}^{N_{P}-1}_{1}$ is the finite horizon of load ports with $N_{P}-1$ being the last port, and $\gamma$ is the discounting factor.
State. The state is given by ${s}_{p}\in{S}$, defined as ${s}_{p}=(u_{p},q_{p},\zeta)$. This includes vessel utilization $u_{p}\in\mathbb{R}_{\geq 0}^{n_{u}}$ and realized demand $q_{p}\in\mathbb{R}_{\geq 0}^{n_{q}}$, where $n_{u}=|B|\times|D|\times|K|\times|\textit{TR}|$, $n_{c}=|B|\times|D|$ and $n_{q}=|K|\times|\textit{TR}|$ are the elements in utilization, location and demand, respectively.
Additionally, environment parameters $\zeta$ contain the expected value $\mu\in\mathbb{R}^{n_{q}}_{\geq 0}$ and standard deviation $\sigma\in\mathbb{R}^{n_{q}}_{>0}$ of demand, load ports $i$, discharge ports $j$ and cargo types $k$ as $(i,j,k)\in\textit{TR}\times K$, TEU per container $\textit{teu}\in\{1,2\}^{n_{q}}$, cargo weight $w\in\mathbb{R}^{n_{q}}_{>0}$, cargo revenue $\textit{rev}\in\mathbb{R}^{n_{q}}_{>0}$.
The initial state $s_{0}=(u_{0},q_{0},\zeta)$ consists of an empty vessel $u_{0}=\mathbf{0}^{n_{u}}$, realized demand $q_{0}$ at initial port, and $\zeta$ is initialized randomly for each episode.
Action. An action ${x}_{p}\in{X}$ assigns a real number of containers to utilization $u_{p}$, where ${x}_{p}\in\mathbb{R}_{>0}^{n_{u}}$ and thus similar to $u^{+}_{p}$. Each action ${x}_{p}$ is subject to a feasible region, defined by polyhedron $\textit{PH}({s}_{p})=\{{x}_{p}\in\mathbb{R}^{n_{u}}_{>0}:A({s}_{p}){x}_{p}\leq b({s}_{p})\}$. Here, $A({s}_{p})\in\mathbb{R}^{m_{u}\times{n_{u}}}$ is the constraint matrix, $b({s}_{p})\in\mathbb{R}^{m_{u}}$ is the bound vector, and $m_{u}$ is the number of constraints.
Previous work demonstrated that linearly relaxed MPPs outperform exact MPPs, as the subsequent SPP effectively discretizes the solution Pacino et al. (2011). Accordingly, we can adopt real-valued actions instead of integer-valued actions. Given a traditional MPP formulation with utilization $u_{p}$ Van Twiller et al. (2024a), we define constraints of $\textit{PH}({s}_{p})$ in terms of actions ${x}_{p}$ and pre-loading utilization $u^{\prime}_{p}$.
$\quad{x}_{p}^{\top}\mathbf{1}_{n_{c}}$ $\leq\;q_{p}$ (3)
$\quad{x}_{p}\textit{teu}$ $\leq\;c-u^{\prime}_{p}\textit{teu}$ (4)
-$\text{-}\mathbf{1}^{\top}\!\big{(}(\textit{lm}\!-\!\underline{\textit{lcg}}w)\!\odot\!{x}_{p}\big{)}$$\leq\!\text{-}\mathbf{1}^{\top}\!\big{(}(\underline{\textit{lcg}}w\!-\!\textit{lm})\!\odot\!u^{\prime}_{p}\big{)}$ (5)
 $\;\mathbf{1}^{\top}\!\big{(}(\textit{lm}\!-\!\overline{\textit{lcg}}w)\!\odot\!{x}_{p}\big{)}$  $\leq\!\;\mathbf{1}^{\top}\!\big{(}(\overline{\textit{lcg}}w\!-\!\textit{lm})\!\odot\!u^{\prime}_{p}\big{)}$ (6)
-$\text{-}\mathbf{1}^{\top}\!\big{(}(\textit{vm}\!-\!\underline{\textit{vcg}}w)\!\odot\!{x}_{p}\big{)}$$\leq\!\text{-}\mathbf{1}^{\top}\!\big{(}(\underline{\textit{vcg}}w\!-\!\textit{vm})\!\odot\!u^{\prime}_{p}\big{)}$ (7)
 $\;\mathbf{1}^{\top}\!\big{(}(\textit{vm}\!-\!\overline{\textit{vcg}}w)\!\odot\!{x}_{p}\big{)}$  $\leq\!\;\mathbf{1}^{\top}\!\big{(}(\overline{\textit{vcg}}w\!-\!\textit{vm})\!\odot\!u^{\prime}_{p}\big{)}$ (8)
Constraint (3) limits load ${x}_{p}$ to the available demand $q_{p}$, while Constraint (4) ensures ${x}_{p}$ does not exceed the residual TEU capacity. Constraints (5)–(6) enforce lcg limits, while Constraints(7)–(8) impose similar bounds on the vcg. These stability bounds are derived from Constraints (1) and (2).
Transition. We use a stochastic transition function $\mathcal{T}({s}_{p+1}|{s}_{p},{x}_{p},)\in\Delta({S})$. Within an episode, the transition consists of multiple components:
  • Upon port arrival, port demand $q_{p}$ is revealed.
  • Subsequently, onboard cargo is discharged $u_{p+1}=u_{p}\odot(1-\mathbf{e}_{p}^{-})$, where $\mathbf{e}_{p}^{-}\in\{0,1\}^{n_{q}}$ is a binary mask indicating the cargo type and transport to nullify in $u_{p}$.
  • Finally, cargo is loaded onboard $u_{p+1}=u_{p}+{x}_{p}$. Action ${x}_{p}$ is based on the current utilization $u_{p}$ and revealed demand $q_{p}$ of port $p$. Future port demand stays unknown.
Reward. Equation (4.1) defines our deterministic reward function, computing profit as the difference between revenue and costs. Revenue is computed as the sum of $x_{p}$ corresponding to elements of $q_{p}$. Since revenue cannot be obtained from containers exceeding demand, the summation of $x_{p}$ is restricted to the elements within $q_{p}$. While costs are computed via the state-dependent auxiliary variables hatch overstows $\textit{ho}({s}_{p},p)\in\mathbb{R}^{|B|}_{>0}$ and excess crane moves $\textit{cm}({s}_{p},p)\in\mathbb{R}^{|B|-1}_{>0}$. These costs are weighted by $\textit{ct}^{\textit{ho}}\in\mathbb{R}_{>0}$ and $\textit{ct}^{\textit{cm}}\in\mathbb{R}_{>0}$, respectively.
$\mathcal{R}(s,x,p)=$ $\;\textit{rev}\min\big{(}x^{\top}\mathbf{1}_{n_{c}},q\big{)}$
$-\textit{ct}^{\textit{ho}}\mathbf{1}^{\top}\textit{ho}(s,p)+\textit{ct}^{\textit{cm}}\mathbf{1}^{\top}\textit{cm}(s,p).$ (9)
4.2 Decomposed MDP
The formal MDP defines an effective action space of size $|{X}|\propto|B|\cdot|D|\cdot|K|\cdot|P|$, where each action ${x}_{p}$ determines how cargo types and transport options are placed on the vessel per port. However, this action space is large, which can hinder learning efficiency Kanervisto et al. (2020).
To address this, we decompose the formal MDP into granular, sequential steps based on an index $(i,j,k)\;\forall(i,j)\in\textit{TR},k\in K$. Instead of placing all transport and cargo types simultaneously, we take a decomposed action for all transports $(p,j)$ and cargo types $k$ at port $p$, then departing to a new port. This reduces the action space to $|X|=|B|\cdot|D|$, while unfolding transports and cargo types over an extended time horizon $t\in H=\{0,1,\dots,T_{\textit{seq}}\}$ with $T_{\textit{seq}}=|K|\cdot|\textit{TR}|.$
State. The state $s_{t}=(u_{t},q_{t},\zeta)$ depends on time step $t$, where $u_{t}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is vessel utilization, and $q_{t}\in\mathbb{R}^{n_{q}}_{\geq 0}$ is realized demand. The environment parameter $\zeta$ remains unchanged. Given the time $t$, however, we can extract relevant parameters from $\zeta$, such as $(\textit{pol}_{t},\textit{pol}_{t},k_{t}),\textit{rev}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$.
Action. Action $x_{t}\in\mathbb{R}^{n_{c}}_{\geq 0}$ assigns real number of containers to utilization $u_{t}$ for step $t$. Each action is subject to $\textit{PH}(s_{t})=\{x_{t}\in\mathbb{R}^{n_{c}}_{>0}:A(s_{t})x_{t}\leq b(s_{t})\}$. Here, $A(s_{t})\in\mathbb{R}^{m_{c}\times{n_{c}}}$ is the constraint matrix, $b(s_{t})\in\mathbb{R}^{m_{c}}$ is the bound vector, and $m_{c}$ is the number of constraints. It is also worth noting that Constraints (3)-(8) are reformulated to fit feasible region $\textit{PH}(s_{t})$.
Transition. At each time step $t$, the transition includes loading, where $x_{t}$ is added to $u_{t}$. However, discharging and demand realization occur only when arriving at a new port, indicated by $t\in T_{\text{new port}}$.
Reward. The revenue at step $t$ is computed as $\textit{rev}(\textit{pol}_{t},\textit{pod}_{t},k_{t})\min(\mathbf{1}^{\top}x_{t},q_{t}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}).$ However, costs depend on knowing all loading operations at port $p$, which is aggregated in utilization $u_{t}$ at the last step of the port $t\in T_{\text{leave port}}$. As a result, the cost signal is sparse, being evaluated only once per port $p$, rather than at each step.
5 Proposed Architecture
Our approach consists of several components: an encoder-decoder model parameterized by $\theta$, an actor-critic DRL method, and a feasibility projection layer. The encoder-decoder model shown in Figure 2 employs a look-ahead policy $\pi_{\theta}(x|s_{t})$ conditioned on state $s_{t}$ and parameterized by mean $\mu_{\theta}(s_{t})$ and standard deviation $\sigma_{\theta}(s_{t})$. By training in the decomposed MDP and optionally projecting the policy, we iteratively generate actions $x_{t}$ to construct solutions.
Figure 2: Deep Reinforcement Learning Architecture with Feasibility Projection for Actor-Critic Methods
5.1 Encoder-Decoder Model
Figure 3 presents our encoder-decoder model, based on that of Kool et al. (2019), with modifications highlighted below.
Figure 3: Layers of the encoder and the actor-critic decoder.
Cargo Embedding. The cargo embedding $e^{\text{cargo}}(\zeta)$ parameterized by $\theta_{\textit{cr}}$ maps cargo-related episode information in $\zeta$ into a feature representation for the encoder. To enhance positional awareness, we subsequently apply sinusoidal positional encoding Vaswani et al. (2017).
Encoding Layers. An attention encoder $f(e^{\text{cargo}}(\zeta))$ parameterized by ${\theta_{\textit{enc}}}$ maps embedding $e^{\text{cargo}}(\zeta)$ to latent variable $z$ using multi-head attention (MHA) to identify revalant features dynamically Vaswani et al. (2017). Then, we use a feed-forward network (FFN) with ReLU activation, layer normalization, residual connections, and dropout.
Context Embedding. The context embedding $e^{\text{cont}}(u_{t},z)$, parameterized by ${\theta_{\textit{co}}}$, focuses on time $t$, extracting time-specific features from the utilization $u_{t}$ and latent variable $z$. These features provide the MHA query with a representation of the vessel’s current condition, enabling the policy to make decisions based on the present context.
Dynamic Embedding. The dynamic embedding $e^{\text{dynam}}(q_{t},z)$, parameterized by ${\theta_{\textit{dn}}}$, extracts features from demand $q_{t}$ in all steps of horizon $H$, capturing patterns across the entire episode. It combines real demand and latent information to produce keys and values for the MHA layer. As such, it accounts for global temporal trends, allowing it to anticipate future conditions and make long-term decisions.
Actor Layers. As proposed in Kool et al. (2019), our actor decoder $g(e^{\text{cont}}(u_{t},z),e^{\text{dynam}}(q_{t},z))$ is an attention model parameterized by $\theta_{\textit{act}}$. However, our actor takes input from the context and dynamic embedding. An MHA layer with a forward-looking mask bases decisions on steps ${t,t+1,\dots,T_{\textit{seq}}}$ to anticipate future events. Subsequently, an FFN extracts features, after which a pointer mechanism performs a soft selection of relevant steps using key-value inputs Vinyals et al. (2015). A softplus activation outputs positive logits $({\mu}_{\theta}(s_{t}),{\sigma}_{\theta}(s_{t}))$.
Critic Layers. Our critic model $V(s_{t},z)$, parameterized by $\theta_{\textit{cri}}$, estimates the value of state $s_{t}$ and latent variable $z$ through an FFN outputting $V_{\theta_{\textit{cri}}}(s_{t},z)\in\mathbb{R}$.
Action Policy. The actor logits parameterize a stochastic policy $\pi_{\theta}(x|s_{t})=\mathcal{N}(({\mu}_{\theta}(s_{t}),{\sigma}_{\theta}(s_{t}))$, which allows action sampling $x_{t}\sim\pi_{\theta}(x|s_{t})$ to construct solutions.
5.2 Feasibility Regularization in Actor-Critic Loss
Standard actor-critic methods can efficiently and stably learn policies that require implicit feasibility in the action space Schulman et al. (2017); Haarnoja et al. (2018), while explicit constraints are often integrated by feasibility regularization (FR) in the actor loss Ding et al. (2020); Calvo-Fullana et al. (2021). Equation (11) defines a composite loss that combines the actor loss $\mathcal{L}_{\text{actor}}(\theta)$ with a soft regularization term $\mathcal{L}_{\text{feas}}(\theta)$. FR is defined in Equation (11), where $\lambda_{f}$ controls the regularization strength, and policy samples $x_{\theta}(s_{t})$ are assumed to allow gradient flow to update parameters $\theta$, as is the case in SAC Haarnoja et al. (2018). For algorithms that do not include actions in the computation graph, e.g., PPO Schulman et al. (2017), we substitute $x_{\theta}(s_{t})$ with $\mu_{\theta}(s_{t})$.
Determining Lagrangian multipliers ($\lambda_{f}$) for multiple constraints with varying scales is already challenging for static feasible regions. It requires costly hyperparameter tuning to balance a trade-off between objectives and feasibility effectively. As a result, FR is a naive approach when applied to dynamic, state-dependent feasible regions $\textit{PH}(s_{t})$ Calvo-Fullana et al. (2021); Donti et al. (2021). However, FR remains useful as a comparative baseline.
$\mathcal{L}(\theta)=$ $-\mathcal{L}_{\text{actor}}(\theta)+\lambda_{{f}}\mathcal{L}_{\text{feas}}(\theta)$ (10)
$\mathcal{L}^{\text{feas}}(\theta)=$ $\;\mathbb{E}_{t}\left[\left(\max(A(s_{t}){x}_{\theta}(s_{t})-b(s_{t}))_{>0}\right)^{2}\right]$ (11)
5.3 Feasibility Layers
To extend beyond FR, we leverage differentiable projection layers for constraint minimization. We incorporate two constraint-specific and one general layer(s). Note that a single action may be infeasible while combining such actions can achieve feasibility. Therefore, we prioritize violation minimization over strict enforcement to allow for flexibility.
Furthermore, non-linear transformations of continuous distribution samples change their probability density. We account for this transformation using Jacobian adjustments to the policy’s log probabilities Bishop (2006).
Technical details on feasibility layers are provided in Appendix B of supplementary material.
Weighted Scaling Projection. Function (12) defines the weighted scaling layer (WS), which normalizes vector $x$ if the sum of $x$ exceeds scalar $y$. This preserves relative proportions of elements in $x$ while enforcing $x$ to sum to scalar $y$.
$\mathcal{W}(\mathbf{x},y)=\begin{cases}y\frac{\mathbf{x}}{\mathbf{1}^{\top}\mathbf{x}}&\text{if }\mathbf{1}^{\top}\mathbf{x}>y\\ \mathbf{x}&\text{otherwise}\end{cases}$ (12)
Policy Clipping. We can apply function $\mathcal{C}(x,\textit{lb}_{\textit{pc}},\textit{ub}_{\textit{pc}})=\max(\min(x,\textit{ub}_{\textit{pc}}),\textit{lb}_{\textit{pc}})$ to enforce element-wise limits on vector $x$, thereby performing policy clipping (PC) on actions. However, PC is only applicable to box constraints.
Violation Projection. In Algorithm 1, we define a violation projection (VP) layer that reduces inequality constraint violations by shifting point $x$ closer to the feasible region of some convex polyhedron $\textit{PH}=\{x\in\mathbb{R}^{n}_{>0}:Ax\leq b\}$, where $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$. The inequality $Ax\leq b$ ensures PH is convex, enabling gradient-based reduction of violation $\mathcal{V}(x)$ Boyd and Vandenberghe (2004). To measure feasibility, we compute the element-wise violation $\mathcal{V}(x)=(Ax-b)_{>0}$, where each $\mathcal{V}(x)_{m_{i}}>0$ indicates the violation of constraint ${}_{m_{i}}$, and $\mathcal{V}(x)_{m_{i}}=0$ indicates satisfaction of that constraint. To minimize violations, we iteratively update $x$ by applying gradient descent to reduce the violation term $\|\mathcal{V}(x)\|_{2}^{2}$, representing the squared distance from $x$ to PH. Equation (13) updates $x$ with a step size $\eta_{v}$, which is rewritten to the update function found in Algorithm 1.
$x^{\prime}$ $=x-\eta_{v}\nabla_{x}\|\mathcal{V}(x)\|_{2}^{2}$ (13)
During training, the VP layer continues for a fixed number of epochs. However, we include a stopping criteria during inference, which continues until the change in total violation $\mathbf{1}^{\top}\mathcal{V}(x^{\prime})-\mathbf{1}^{\top}\mathcal{V}(x)$ is below a threshold $\delta_{v}$. As a result, we $x$’s distance to the feasible region and incorporate constraint awareness into initially unconstrained policies.
Algorithm 1 Violation Projection Layer
0:  $x\in\mathbb{R}^{n}_{>0}$, parameters $(A,b,\eta_{v},\delta_{v})$
1:  Initialize $x^{\prime}\leftarrow x$
2:  Define $\mathcal{V}(x)\leftarrow(Ax-b)_{>0}$
3:  for $i=1$ to epochs do
4:     Set $x\leftarrow x^{\prime}$
5:     Update $x^{\prime}\leftarrow x-\eta_{v}A^{\top}\mathcal{V}(x)$
6:     if $\mathbf{1}^{\top}\mathcal{V}(x^{\prime})-\mathbf{1}^{\top}\mathcal{V}(x)\leq\delta_{v}$ then
7:        break
8:     end if
9:  end for
10:  return  $x^{\prime}$
Methods Testing ($\boldsymbol{N=30}$) Generalization ($\boldsymbol{N=30}$)
Alg. Model F.M. Ob. (\$) Time (s) F.I. (%) F.O. (\$) Ob. (\$) Time (s) F.I.(%) F.O. (\$)
SAC AM FR 1113.03 12.63 0.00 1065.80 1211.87 15.35 0.00 1041.42
SAC AM-P FR/VP 1318.05 15.20 100.00 - 1288.68 14.18 100.00 -
SAC AM-P VP 1447.69 15.03 100.00 - 1411.34 13.33 100.00 -
SAC AM-P FR/WS/PC 1373.24 13.34 100.00 - 1340.12 14.42 100.00 -
SAC AM-P WS/PC 1494.22 13.12 100.00 - 1482.40 12.91 100.00 -
PPO AM FR 1842.46 11.74 0.00 1063.24 1830.95 12.60 0.00 1066.69
PPO AM-P FR/VP 1355.45 15.60 100.00 - 1321.90 14.91 100.00 -
PPO AM-P VP 1318.10 14.51 100.00 - 1282.10 14.64 100.00 -
PPO AM-P FR/WS/PC 1369.03 14.29 100.00 - 1330.10 14.17 100.00 -
PPO AM-P WS/PC 1471.40 13.75 100.00 - 1455.98 13.21 100.00 -
CPL SMIP-NA - 1053.02 8434.58 100.00 - 1023.04 8434.94 100.00 -
CPL SMIP-PI* - 1713.73 40.75 100.00 - 1680.51 38.93 100.00 -
Table 1: Experimental results comparing DRL methods (SAC/PPO) with a vanilla (AM) or projected attention model (AM-P) and feasibility mechanisms (F.M.): feasibility regularization (FR), weighted scaling (WS), policy clipping (PC) and violation projection (VP). We compare SMIP-NA (non-anticipation) as a baseline, and SMIP-PI (* assumes perfect information which is unrealistic) as an expected upper bound, both solved with CPLEX (CPL). Average performance metrics on $N$ instances include objective value in profit (Ob.), inference time in seconds (Time), percentage of feasible instances (F.I.), and objective value with feasibility recovery (F.O.). Note that † indicates infeasible objectives. Generalization performance is evaluated on unseen, out-of-distribution instances based on a uniform distribution.
6 Experimental Results
We compare our policies against baselines on in-distribution and out-of-distribution instances, evaluating the objective value, computational cost and feasibility. Additionally, an ablation study analyzes the impact of feasibility mechanisms and gradient flow from actions, while managerial insights discuss the value of information, computational cost, and adaptiveness to varying levels of uncertainty.
6.1 Experimental Setup
Instance Generation. Training instances are sampled from a Gaussian distribution $\mathcal{N}(\mu^{(i,j,k)},\sigma^{(i,j,k)})\;\forall(i,j)\in\textit{TR},k\in K$, where $\mu^{(i,j,k)}$ is randomly generated and $\sigma^{(i,j,k)}=\textit{CV}\mu^{(i,j,k)}$ controls spread via the coefficient of variation CV. A continuous uniform distribution generates out-of-distribution instances $\mathcal{U}(\textit{lb}^{(i,j,k)},\textit{ub}^{(i,j,k)})\;\forall(i,j)\in\textit{TR},k\in K$ to test generalization. Details are in Appendix C.
Feasibility Implementation. We evaluate feasibility by integrating multiple mechanisms to address all constraints in $\textit{PH}(s_{t})$. Starting with a baseline using FR, we introduce two alternatives: a constraint-specific (FR/WS/PC) and a general projection approach (FR/VP). We then remove FR from both to assess its impact. WS and PC are not analyzed separately, as it leaves many constraints unaddressed.
SMIP Baseline. We compare our approach against two common techniques in stochastic programming: SMIP-NA (stochastic MIP without anticipation) and SMIP-PI (stochastic MIP with perfect foresight, serving as an upper bound). The scenario trees and MIP models are shown in Appendix D.
Runs. Training occurs offline on simulated instances for a 1,000 TEU vessel over a 4-port voyage. GPU-based experiments use an NVIDIA RTX A6000, and CPU-based runs use an AMD EPYC 9454 48-core processor. Implementation details and additional experiments are provided in Appendix E.
6.2 Policy Performance
Table 1 compares the performance of different solution approaches. Our projected attention models (AM-P) outperform SMIP-NA, achieving around 30-35% higher objective values and 500× faster computations. The AM with FR misleadingly attains higher profits than SMIP-NA due to infeasibility. If we recover feasibility, then profits are significantly reduced to the level of SMIP-NA. In contrast, each AM-P finds feasible solutions with a higher profit than AM with FR, highlighting FR’s limitations. All AM-P policies generalize well, with SAC and PPO showing an average profit reduction of 2% and 3% on unseen uniform instances, similar to SMIP-NA and SMIP-PI. These results suggest that AM-P can leverage imperfect information to anticipate future uncertainty and ensure feasibility by projection layers beyond FR.
Ablation Study. Table 1 analyzes the impact of feasibility mechanisms. Removing FR from FR/VP increases profit for SAC and slightly decreases profit for PPO, without diminishing feasibility. When FR is removed from FR/WS/PC, profit increases, and both DRL approaches continue to maintain feasibility. This outcome is surprising, given the absence of a direct mechanism to account for stability. However, from our domain understanding, we know container vessels generally become more stable as utilization increases in all locations. Replacing SAC with PPO, which eliminates action gradient flow, reduces feasibility in FR, though projection layers mitigate this effect. These findings highlight the impact of feasibility mechanisms on the objective value and feasibility.
6.3 Managerial Insights
Figure 4 performs a sensitivity analysis to examine the effects of SMIP scenario size and demand uncertainty, reporting with an average and 95% confidence interval (CI).
Value of Information. Figure 4(a) compares performance under non-anticipation, imperfect, and perfect information. The predictive accuracy of SMIP models improves with the number of scenarios, as shown by the stabilizing yet decreasing objective values. With 28 scenarios, perfect information enhances profit by 60-65%, while imperfect information via DRL with VP achieves a 25-40% improvement, emphasizing the role of information availability for decision quality.
Computational Time. While inference time is critical, training time remains relevant. Figure 4(b) illustrates the total computational time of 30 instances, with training time included for DRL. Computational cost increases significantly with the number of scenarios, with SMIP-NA exhibiting exponential growth and SMIP-PI showing steady growth. Solving the most accurate SMIP-NA takes approximately 2 hours per instance, which is intractable for stowage planners. In contrast, the DRL approach requires offline training of about 11 hours, after which solutions can be constructed in seconds.
Adaptiveness. Figure 4(c) illustrates how policies adapt to variations in the spread of the demand distribution of unseen instances. As variability and uncertainty increase, both SAC and PPO policies achieve higher profits, demonstrating their ability to leverage uncertainty. Despite fluctuations in variability, both policies consistently generate feasible solutions.
Implications for Practical Deployment. Our findings are based on simulations, limited by the lack of real-world data. To ensure effective deployment, practitioners should focus on collecting and generating representative cargo demand data.
4444888812121212161616162020202024242424282828281,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 800Number of ScenariosProfit (\$)SMIP-NASMIP-PI PPO/VPSAC/VP
(a) Profit with 95% CI across scenario sizes on 30 instances.
[…]
(b) Total computational time across scenario sizes of 30 instances.
4444888812121212161616162020202024242424282828281,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8000.10.10.10.10.30.30.30.30.50.50.50.50.70.70.70.70.90.90.90.91,30013001{,}3001 , 3001,40014001{,}4001 , 4001,50015001{,}5001 , 500Coefficient of Variation (CV)Profit (\$)PPO/VPSAC/VPSMIP-NASMIP-PI PPO/VPSAC/VP
(c) Profit with 95% CI across CV levels on 30 unseen instances.
Figure 4: Sensitivity analysis of scenario size and demand spread.
7 Conclusion and Future Directions
This work introduces a novel MDP formulation for the MPP under demand uncertainty, incorporating realistic inequality constraints. We train an AM policy using actor-critic DRL methods with differentiable feasibility projections to construct MPP solutions. Experimental results demonstrate that our policy efficiently generates adaptive and feasible solutions, significantly outperforming baseline DRL methods and the SMIP-NA. This approach establishes an AI-driven decision-support policy for planning under uncertainty in a critical part of the global supply chain. Future work will extend the MPP formulation and scale to larger vessels and longer voyages, further enhancing the representativeness.
Ethical Statement
Human oversight is critical in decision-support tools to ensure fairness, accountability, and transparency. While models can enhance decision-making, they must not replace human judgment, particularly in high-stakes applications. We strongly encourage safeguards to mitigate automation bias, ensuring that users can override or challenge outputs, with continuous validation and auditing to uphold trust and fairness.
Appendix A MDP of Master Planning Problem
A.1 Sets and Parameters
Provided the sets and parameters in Section 2, we introduce additional subsets of the transport set TR given port $p\in{P}$:
  • Onboard transports: $\textit{TR}^{\textit{OB}}_{p}=\{(i,j)\!\in\!{P}^{2}\!\mid\!i\leq p,j>p\}$
  • Arrival transports: $\textit{TR}^{\textit{AC}}_{p}=\{(i,j)\!\in\!{P}^{2}\!\mid\!i<p,j>p\}$
  • Load transports: $\textit{TR}^{{+}}_{p}=\{(p,j)\!\in\!{P}^{2}\!\mid\!j>p\}$
  • Discharge transports: $\textit{TR}^{{-}}_{p}=\{(i,p)\!\in\!{P}^{2}\!\mid\!i<p\}\;$
  • Transports in crane operations: $\textit{TR}^{M}_{p}=\textit{TR}^{+}_{p}\cup\textit{TR}^{-}_{p}$
Considering episode parameters $\zeta$, we define:
  • Transports: $\textit{tr}=(i,j)\in\textit{TR}$
  • Cargo types: $k=(\kappa_{1},\kappa_{2},\kappa_{3})\in K$
For each combination $(i,j,k)$, we associate the expected demand $\mu^{(i,j,k)}$, standard deviation $\sigma^{(i,j,k)}$, TEU per container $\textit{teu}^{(i,j,k)}$, container weight $w^{(i,j,k)}$, and revenue per container $\textit{rev}^{(i,j,k)}$.
The TEU per container depends on $k$ as:
$\textit{teu}(k)=\begin{cases}1,&\text{if }\kappa_{1}=\text{20 ft.}\\ 2,&\text{if }\kappa_{1}=\text{40 ft.}\end{cases}\;.$
Similarly, the container weight is defined by:
$\textit{w}(k)=\begin{cases}1,&\text{if }\kappa_{2}=\text{Light}\\ 2,&\text{if }\kappa_{2}=\text{Medium}\\ 3,&\text{if }\kappa_{2}=\text{Heavy}\end{cases}\;.$
The revenue function is given by:
$\textit{rev}(i,j,k)=\begin{cases}(j-i)(1-\textit{LR})\!+\!0.1,&\text{if }\kappa_{3}=\text{Long}\\ (j-i)\!+\!0.1,&\text{if }\kappa_{3}=\text{Spot}\end{cases}\;.$
The parameters $\mu$ and $\sigma$ are randomly generated, as shown in Appendix C.
A.2 Formal MDP
We define the MDP by decomposing the traditional CO problem outlined in Van Twiller et al. [2024a].
In the traditional MPP, cargo is loaded onto a vessel at each port in a voyage. Let $u\in\mathbb{R}^{n_{u}}$ represent the vessel’s utilization over the voyage, which is defined by the set of ports $P=\{1,2,\dots,N_{P}\}$. Let us also recall the following:
$n_{u}=|B|\times|D|\times|K|\times|\textit{TR}|$
$n_{c}=|B|\times|D|,\quad n_{q}=|K|\times|\textit{TR}|\quad n_{u}=\{n_{c}\times n_{q}\}$
Utilization $u$ can be decomposed into individual voyage legs, corresponding to the segments between consecutive ports. Specifically, we decompose as:
$u=(u_{0},u_{1},\dots,u_{N_{P-1}})$
where each $u_{p}\in\mathbb{R}^{n_{u}}$ represents the vessel’s utilization immediately after operations at port $p$.
A.2.1 Feasible Region
Suppose we have a feasible region for the MPP, where:
  • $A^{\prime}\in\mathbb{R}^{m_{u}\times n_{u}}$ is the constraint matrix,
  • $b^{\prime}\in\mathbb{R}^{m_{u}}$ is the bound vector,
  • $u_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is the nonnegative vessel utilization.
The feasible region, denoted as PH, is given by:
$\textit{PH}({s}_{p})=\{u_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}\mid A^{\prime}u_{p}\leq b^{\prime}\}\;.$
At port $p$, utilization can be decomposed into load operations and pre-load utilization:
$u_{p}=u^{\prime}_{p}+u^{+}_{p}\;,$
where:
  • $u^{+}_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}$ represents load operations,
  • $u^{\prime}_{p}=u_{p-1}-u^{-}_{p}$ is the utilization before load operations,
  • $u_{p-1}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is the previous step’s utilization,
  • $u^{-}_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is the discharge operations.
Consequently, we can rewrite the feasible region as:
$\textit{PH}({s}_{p})=\{u^{+}_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}\mid A^{\prime}u^{+}_{p}\leq b^{\prime}-A^{\prime}u^{\prime}_{p}\}\;.$
A.3 Decomposed MDP
Utilization can be decomposed into sequential steps to refine temporal granularity, thereby obtaining an decomposed MDP formulation. We decompose $u$ as:
$u=(u_{0},u_{1},\dots,u_{T_{\textit{seq}}})\;,$
where $u_{t}\in\mathbb{R}^{n_{u}}$ represents the utilization at time step $t$, and $t\in H=\{1,2,\dots,T_{\textit{seq}}\}$ denotes the episodic horizon $H$.
Each step $t$ represents a transport and cargo type, as tuple $(\textit{pol}_{t},\textit{pod}_{t},k_{t})$. Algorithm 2 illustrates an episode of the decomposed MDP. First, we reset the state $s_{0}$, initialize time $t$, and an empty trajectory. The episode iterates over load ports ($\textit{pol}_{t}$), discharge ports ($\textit{pod}_{t}$), and cargo classes ($k_{t}$). At each step, we sample action $x_{t}$ from policy $\pi_{\theta}(x|s_{t})$ conditioned on state $s_{t}$ and episode parameters $\zeta$, and transition to state $s_{t+1}$. Afterwards, we store the results in the trajectory and increment time $t$. This process continues until all combinations of $\textit{pol}_{t}$, $\textit{pod}_{t}$, and $k_{t}$ are explored, accumulating a total of $T_{\textit{seq}}$ steps.
Algorithm 2 Episode of Augmented MDP
0:  $\mathcal{T}$, $\pi_{\theta}$, $\zeta$, $\mathcal{Q}$
1:  $q_{T_{\textit{seq}}}^{(i,j,k)}\sim\mathcal{Q}(\mu^{(i,j,k)},\sigma^{(i,j,k)})\;\forall(i,j)\in\textit{TR},k\in K$
[…]
3:  for $\textit{pol}_{t}=1$ to $N_{P}-1$ do
4:     for $\textit{pod}_{t}=\textit{pol}_{t}+1$ to $N_{P}$ do
5:        for $k_{t}\in K$ do
6:           $x_{t}\sim\pi_{\theta}(x|s_{t},\zeta)$
7:           $s_{t+1}\sim\mathcal{T}(s_{t},x_{t},\zeta)$
8:           Append $(s_{t},x_{t},r_{t},s_{t+1})$ to Trajectory
9:           $t\leftarrow t+1$
10:        end for
11:     end for
12:  end for
13:  return  Trajectory
A.3.1 Transitions.
We use a stochastic transition function $\mathcal{T}(s_{t+1}|s_{t},x_{t},\zeta)\in\Delta(S)$. The transition consists of sequential steps:
  1. 1.
    If $t\in T_{\text{new port}}$, port demand is revealed. This means we show $q_{t}^{(i,j,k)}\;\forall(i,j)\in\textit{TR}^{+}_{\textit{pol}_{t}},k\in K$.
  2. 2.
    If $t\in T_{\text{new port}}$, onboard cargo is discharged $u_{t+1}=u_{t}\odot(1-\mathbf{e}_{t}^{-})$, where $\mathbf{e}_{t}^{-}\in\{0,1\}^{n_{q}}$ is a binary mask indicating the cargo type and transport to nullify in $u_{t}$.
  3. 3.
    Each time $t$, cargo is loaded onboard $u_{t+1}=u_{t}+x_{t}\odot\mathbf{e}_{t}^{+}$, where $\mathbf{e}_{t}^{+}\in\{0,1\}^{n_{q}}$ is a binary indicator specifying cargo types and transports to add to $u_{t}$.
Additionally, we define the set of time steps before we leave for a new port $p+1$ is defined as follows:
$T_{\text{leave port}}$ =$=\Big{\{}t\in H\ \mid\exists p\in P^{N_{P}-1}_{1}\text{ such that }$
$\qquad\qquad t=|K|\Bigl{(}p(N_{P}\!-1\!)-\frac{p(p-1)}{2}\Bigr{)}-1\Big{\}}.$
Finally, the set of time steps at which we arrive at a new port $p$ is defined as follows:
$T_{\text{new port}}$ =$=\Big{\{}t\in H\ \mid\exists p\in P^{N_{P}-1}_{1}\text{ such that }$
$\qquad\qquad t=|K|\Bigl{(}(p-1)(N_{P}\!-1\!)-\frac{p(p-1)}{2}\Bigr{)}\Big{\}}.$
A.3.2 Feasible Region
The state-dependent feasible region for each time $t$ is formulated as:
$\textit{PH}(s_{t})=\{u_{t}\in\mathbb{R}^{n_{u}}_{\geq 0}\mid A^{\prime}u_{t}\leq b^{\prime}\}$
Similar to the port utilization, utilization can be decomposed into load operations and pre-load utilization:
$u_{t}=u^{\prime}_{t}+u^{+}_{t}$
where:
  • $u^{+}_{t}\in\mathbb{R}^{n_{u}}_{\geq 0}$ represents load operations,
  • $u^{\prime}_{t}=u_{t-1}-u^{-}_{t}$ is the utilization before load operations,
  • $u_{t-1}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is the previous step’s utilization,
  • $u^{-}_{t}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is the discharge operations.
Using the decomposition, we obtain the feasible region as:
$\textit{PH}(s_{t})=\{u_{t}\in\mathbb{R}^{n_{u}}_{\geq 0}\mid A^{\prime}(u^{+}_{t}+u^{\prime}_{t})\leq b^{\prime}\}$
A.3.3 Substituting Load Operations for Actions
Actions $x_{t}$ correspond to transformed load operations $u^{+}_{t}$, given by:
$x_{t}=u^{+}_{t}M(s_{t}),\quad M(s_{t})\in\{0,1\}^{n_{u}\times n_{c}},$
where $M(s_{t})$ is a state-dependent sparsity mask that selects relevant elements from $u^{+}_{t}$.
However, load operations are subject to $m_{u}$ constraints, whereas actions adhere to $m_{c}$ constraints. To bridge this difference, we define the state-dependent constraint matrix:
$A(s_{t})=T(s_{t})^{\top}A^{\prime}M(s_{t}),\quad T(s_{t})\in\{0,1\}^{m_{u}\times m_{c}},$
where:
  • $A^{\prime}$ is the original constraint matrix of shape $(m_{u},n_{u})$,
  • $T(s_{t})$ maps the constraints of $u_{t}^{+}$ to that of $x_{t}$
  • $M(s_{t})$ maps the space of $u_{t}^{+}$ to that of $x_{t}$,
Similarly, we introduce a state-dependent bound:
$b^{\prime\prime}(s_{t})=T(s_{t})^{\top}b^{\prime},$
where:
  • $b^{\prime}$ is the original bound of shape $(m_{u},1)$,
  • $T(s_{t})$ maps the constraints of $u_{t}^{+}$ to that of $x_{t}$
A.3.4 Feasible Region for Actions
Using the refined notation, we express the state-dependent feasible region in terms of actions:
$\textit{PH}(s_{t})=\{x_{t}\in\mathbb{R}^{n_{c}}_{\geq 0}\mid A(s_{t})x_{t}\leq b^{\prime\prime}(s_{t})-A^{\prime}u^{\prime}_{t}\}.$
Next, we define the updated bound as:
$b(s_{t})=b^{\prime\prime}(s_{t})-A^{\prime}u^{\prime}_{t}.$
Substituting this into the feasible region, we obtain:
$\textit{PH}(s_{t})=\{x_{t}\in\mathbb{R}^{n_{c}}_{\geq 0}\mid A(s_{t})x_{t}\leq b(s_{t})\}.$
A.4 MPP Constraints
Let us specify the MPP constraints of $\textit{PH}({s}_{p})$ and $\textit{PH}(s_{t})$.
A.4.1 Demand Constraints
Let us consider the demand subset of $\textit{PH}({s}_{p})$ as:
$\textit{PH}({s}_{p})_{\text{dem}}=\{{x}_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}\mid A^{\prime}_{\text{dem}}{x}_{p}\leq b^{\prime}_{\text{dem}}-A^{\prime}_{\text{dem}}u^{\prime}_{p}\}\;.$
We sum over all vessel locations to obtain an aggregated number of containers of shape $n_{q}$. Note that only current load actions ${x}_{p}$ are relevant for $q_{p}$, hence we can omit $A^{\prime}_{\text{dem}}u^{\prime}_{p}$ as pre-loading utilization has already satisfied its demand requirements.
${x}_{p}^{\top}\mathbf{1}_{n_{c}}$ $\leq q_{p}$
Consider the demand subset of $\textit{PH}(s_{t})$ as:
$\textit{PH}(s_{t})_{\text{dem}}=\{x_{t}\in\mathbb{R}^{n_{c}}_{\geq 0}\mid A(s_{t})_{\text{dem}}x_{t}\leq b^{\prime}(s_{t})_{\text{dem}}-A^{\prime}_{\text{dem}}u^{\prime}_{t}\}\;.$
Right now, we can sum the full vector $x_{t}$ as it needs to sum to scalar $q_{t}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$. Again, previous steps are irrelevant to demand, hence we can disregard $A^{\prime}_{\text{dem}}u^{\prime}_{t}$ to obtain:
$\mathbf{1}^{\top}x_{t}$ $\leq q_{t}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$
A.4.2 Capacity Constraints
Let us consider the constraint subset of $\textit{PH}({s}_{p})$ as:
$\textit{PH}({s}_{p})_{\text{cap}}=\{{x}_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}\mid A^{\prime}_{\text{cap}}{x}_{p}\leq b^{\prime}_{\text{cap}}-A^{\prime}_{\text{cap}}u^{\prime}_{p}\}\;.$
We sum TEU of all cargo types and transports in ${x}_{p}$ to obtain TEU use per location with shape $n_{c}$. The TEU of pre-load utilization is also considered by subtracting it from the vessel capacity, obtaining the following:
${x}_{p}\textit{teu}\leq c-u^{\prime}_{p}\textit{teu}\;,$
Consider the demand subset of $\textit{PH}(s_{t})$ as:
$\textit{PH}(s_{t})_{\text{cap}}=\{x_{t}\in\mathbb{R}^{n_{c}}_{\geq 0}\mid A(s_{t})_{\text{cap}}x_{t}\leq b^{\prime}(s_{t})_{\text{cap}}-A^{\prime}_{\text{cap}}u^{\prime}_{t}\}\;.$
Now, we can do the same trick based on a single scalar $\textit{teu}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$ multiplied with the sum of action $x_{t}$.
$^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}\mathbf{1}^{\top}x_{t}\leq c-u^{\prime}_{t}\textit{teu}\;,$ teu ,
A.4.3 Stability Constraints
The stability constraints require some algebra to derive for $\textit{PH}({s}_{p})$ and $\textit{PH}(s_{t})$.
The lcg constraint in its original form is given by:
$\frac{\mathbf{1}^{\top}(\textit{lm}\odot u_{p})}{\mathbf{1}^{\top}(w\odot u_{p})}$ $\leq\overline{\textit{lcg}}\;.$
Applying the utilization decomposition, we can obtain the formulation for $\textit{PH}({s}_{p})$:
$\mathbf{1}^{\top}(\textit{lm}\odot u_{p})$ $\leq\overline{\textit{lcg}}\mathbf{1}^{\top}(w\odot u_{p})$
$\mathbf{1}^{\top}(\textit{lm}\odot u^{+}_{p})+\mathbf{1}^{\top}(\textit{lm}\odot u^{\prime}_{p})$ $\leq\overline{\textit{lcg}}\mathbf{1}^{\top}(w\odot u^{+}_{p})$
$\quad+\overline{\textit{lcg}}\mathbf{1}^{\top}(w\odot u^{\prime}_{p})$
$\mathbf{1}^{\top}(\textit{lm}_{p}\odot{x}_{p})-\overline{\textit{lcg}}\mathbf{1}^{\top}(w\odot{x}_{p})$ $\leq\overline{\textit{lcg}}\mathbf{1}^{\top}(w\odot u^{\prime}_{p})$
$\quad-\mathbf{1}^{\top}(\textit{lm}\odot u^{\prime}_{p})$
 $\;\mathbf{1}^{\top}\!\big{(}(\textit{lm}\!-\!\overline{\textit{lcg}}w)\!\odot\!{x}_{p}\big{)}$  $\leq\!\;\mathbf{1}^{\top}\!\big{(}(\overline{\textit{lcg}}w\!-\!\textit{lm})\!\odot\!u^{\prime}_{p}\big{)}\;.$
This approach extends to both the lower and upper bounds for the lcg and vcg, ensuring that vessel stability is properly maintained at every step.
Based on the $\textit{PH}({s}_{p})$ constraint, we can substitute load operations for actions, and obtain the formulation for $\textit{PH}(s_{t})$:
 $\;\mathbf{1}^{\top}\!\big{(}(\textit{lm}(t)\!-\!\overline{\textit{lcg}}w(t))\!\odot\!x_{t}\big{)}$  $\leq\!\;\mathbf{1}^{\top}\!\big{(}(\overline{\textit{lcg}}w\!-\!\textit{lm})\!\odot\!u^{\prime}_{p}\big{)}\;.$
where $w(t)=w^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$ and $\textit{lm}(t)=\textit{lm}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$
A.5 Auxiliary Variables
The reward function contains two auxiliary variables derived from state $s$, which incur costs due to inefficient port operations. At port $p$, Equation (14) creates an indicator of hatch movements $\textit{hm}(s,p)\in\{0,1\}^{|B|}$, whereas Equation (15) computes the number of on-deck containers during hatch movements, causing hatch overstowage $\textit{ho}(s,p)\in\mathbb{R}^{|B|}_{\geq 0}$.
$\textit{hm}(s,p)$ $=\Bigg{(}\sum_{k\in K}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}u_{t}^{(b,d^{\textit{below}},k,\textit{tr})}>0\Bigg{)}$ (14)
$\textit{ho}(s,p)$ $=\textit{hm}(s,p)\Bigg{(}\sum_{k\in K}\sum_{\textit{tr}\in\textit{TR}^{\textit{AC}}_{p}}u_{t}^{(b,d^{\textit{above}},k,\textit{tr})}\Bigg{)}$ (15)
Equation (16) computes the target crane moves at port $p$ by equally spreading the total demand per port over pairs of adjacent bays, where $\delta^{\textit{cm}}$ is the allowed deviation from the equal spread set by ports. Subsequently, Equation (17) computes the excess crane moves $\textit{cm}(s,p)\in\mathbb{R}^{|B|-1}_{\geq 0}$
$\overline{\textit{cm}}(s,p)$ $=(1+\delta^{\textit{cm}})\frac{2}{|{B}|}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}\sum_{k\in{K}}q^{(\textit{tr},k)}_{t}$ (16)
${\textit{cm}}(s,p)$ $=\max\Bigg{(}\overline{\textit{cm}}(s,p)-\sum_{d\in D}\sum_{k\in K}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}u_{t}^{(0:|B|-1,d,k,\textit{tr})}$
$\qquad\qquad+u_{t}^{(1:|B|,d,k,\textit{tr})},0\Bigg{)}$ (17)
Appendix B Feasibility Mechanisms
Table 2 provides an overview of implemented feasibility mechanisms.
Type Implementation Constraints
FR Composite loss Constraints $\textit{PH}(s_{t})$
VP $\textit{VP}(x_{t},A(s_{t}),b(s_{t}),\alpha_{v},\delta_{v})$ Constraints $\textit{PH}(s_{t})$
WS $\mathcal{W}(x_{t},q_{t})$ Demand $q_{t}$
PC $\mathcal{C}(x_{t},0,c-u^{\prime}_{t}\textit{teu})$ TEU capacity $c$
Table 2: Feasibility mechanisms and relation to constraints
B.1 Log Probability Adjustments
This subsection provides technical details on the adjustments made to the log-probability distribution of the policy as a result of non-linear transformations to distribution samples.
Projecting actions alters the policy’s probability density, necessitating consideration of the change of variables principle Bishop [2006]. This principle ensures valid volume scaling by requiring the transformation $f(x)$ to satisfy:
  1. 1.
    Differentiability: $f(x)$ must be differentiable to compute the Jacobian $J_{f}(x)$ and determine local volume scaling.
  2. 2.
    Non-Singularity: The Jacobian determinant must be non-zero ($\det(J_{f}(x))\neq 0$) to prevent dimensional collapse.
  3. 3.
    Invertibility: $f(x)$ must be locally or globally invertible to ensure a one-to-one mapping between points in the original and transformed spaces.
These properties ensure the transformation is smooth, one-to-one, and well-behaved, enabling the use of the Jacobian adjustment $\log\pi^{\prime}(x|s)=\log\pi(x|s)-\log|\det(J_{f}(x))|$ as a valid probability scaling factor.
B.1.1 Weighted Scaling Projection Layer.
Suppose we have variable $x\in\mathbb{R}^{n}_{>0}$ and scalar $y\in\mathbb{R}_{>0}$ and the following piecewise linear function:
$\mathcal{P}(x,y)=\begin{cases}x&\text{if }\mathbf{1}^{\top}{x}\leq y\\ \frac{x}{\mathbf{1}^{\top}x}\cdot y&\text{if }\mathbf{1}^{\top}{x}>y\end{cases}$
Case 1: $\mathbf{1}^{\top}{x}\leq y$.
$\mathcal{P}(x,y)$ $=x$
$\frac{\partial}{\partial x}\mathcal{P}(x,y)$ $=\frac{\partial x}{\partial x}$
$J_{\mathcal{P}}(x,y)$ $=I_{n}$
Case 2: $\mathbf{1}^{\top}{x}>y$, where we apply the product rule and then the quotient rule of differentiation.
$\mathcal{P}(x,y)$ $=\frac{x}{\mathbf{1}^{\top}x}\cdot y$
$\frac{\partial}{\partial x}\mathcal{P}(x,y)$ $=\frac{\partial}{\partial x}\left(\frac{x}{\mathbf{1}^{\top}x}\cdot y\right)$
$\frac{\partial}{\partial x}\mathcal{P}(x,y)$ $=\frac{\partial}{\partial x}\left(\frac{x}{\mathbf{1}^{\top}x}\right)\cdot y$
$J_{\mathcal{P}}(x,y)$ $=y\cdot\frac{1}{(\mathbf{1}^{\top}x)^{2}}\left(I_{n}\mathbf{1}^{\top}x-x\cdot\mathbf{1}^{\top}\right)$
$J_{\mathcal{P}}(x,y)$ $=\frac{y\cdot I_{n}}{(\mathbf{1}^{\top}x)}-\frac{y\cdot x^{\top}}{(\mathbf{1}^{\top}x)^{2}}$
We obtain the following Jacobian of function $\mathcal{P}(x,y)$:
$J_{\mathcal{P}}(x,y)=\begin{cases}I_{n}&\text{if }\mathbf{1}^{\top}{x}\leq y\\ \frac{y}{(\mathbf{1}^{\top}x)^{2}}\left(I_{n}\mathbf{1}^{\top}x-x1^{\top}\right)&\text{if }\mathbf{1}^{\top}{x}>y\end{cases}$
Finally, we verify that Jacobian adjustment is allowed for the weighted scaling projection by:
  1. 1.
    $\mathcal{P}(x,y)$ has been shown to be differentiable.
  2. 2.
    Provided that $x,y>0$, either case is positive definite as the diagonal elements are strictly positive. We obtain $\det(J_{\mathcal{P}}(x,y))>0$, thus the Jacobian in non-singular.
  3. 3.
    $\mathcal{P}(x,y)$ is locally invertible as $\det(J_{\mathcal{P}}(x,y))\neq 0$.
B.1.2 Violation Projection Layer.
Suppose we have function $\mathcal{P}(x,A,b)=x-\eta_{v}A^{\top}(Ax-b)_{>0}$ with $x\in\mathbb{R}^{n}_{>0}$, $\eta_{v}\in\mathbb{R}_{>0}$, $A\in\mathbb{R}_{\geq 0}^{m\times n}$, $b\in\mathbb{R}^{m}_{\geq 0}$, and $m>n$.
Case 1: $\eta_{v}A^{\top}(Ax-b)=0$.
$\mathcal{P}(x,A,b)$ $=x$
$\frac{\partial}{\partial x}\mathcal{P}(x,A,b)$ $=\frac{\partial x}{\partial x}$
$J_{\mathcal{P}}(x,A,b)$ $=I_{n}$
Case 2: $\eta_{v}A^{\top}(Ax-b)>0$, where we apply the chain rule on the second term.
$\mathcal{P}(x,A,b)$ $=x-\eta_{v}A^{\top}(Ax-b)$
$\frac{\partial}{\partial x}\mathcal{P}(x,A,b)$ $=\frac{\partial}{\partial x}\left(x-\eta_{v}A^{\top}(Ax-b)\right)$
$J_{\mathcal{P}}(x,A,b)$ $=I_{n}-\eta_{v}A^{\top}A$
Both cases are combined in the following matrix formulation with diagonal matrix $\text{Diag}=\text{diag}((Ax-b)>0)$:
$J_{\mathcal{P}}(x,A,b)$ $=I_{n}-\eta_{v}A^{\top}\text{Diag}A$
Finally, we can confirm that Jacobian adjustment is allowed for the violation projection layer by:
  1. 1.
    $\mathcal{P}(x,A,b)$ is differentiable as shown above.
  2. 2.
    Due to the full rank nature of the identity $I_{n}$ and $A^{\top}\text{Diag}A$ when $\text{Diag}=I_{m}$, we get $\det(J_{\mathcal{P}}(x,A,b))\neq 0$, and hence the Jacobian is non-singular.
  3. 3.
    $\mathcal{P}(x,A,b)$ is locally invertible as $\det(J_{\mathcal{P}}(x,A,b))\neq 0$. It is not globally invertible, due to the piece-wise nature of $(Ax-b)_{>0}$.
B.1.3 Policy Clipping
We can implement a clipped Gaussian distribution that enforces element-wise bounds on a standard Gaussian Fujita and Maeda [2018]. Let $\mu_{\theta}$ and $\sigma^{2}_{\theta}$ denote the policy’s mean and variance, with bounds $\textit{lb}_{\textit{pc}}$ and $\textit{ub}_{\textit{pc}}$, and $\Phi(\cdot)$ being the cumulative distribution function of the standard Gaussian. Actions are sampled from $\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta})$ and clipping the result to $[\textit{lb}_{\textit{pc}},\textit{ub}_{\textit{pc}}]$. Provided this transformation, we compute the log probabilities $\log\pi(x|s)$ for action $x$ by:
$\log\pi(x|s)=\begin{cases}\log\Phi\left(\frac{\textit{lb}_{\textit{pc}}-\mu_{\theta}}{\sigma_{\theta}}\right)&\text{if }x\leq\textit{lb}_{\textit{pc}},\\ -\frac{(x-\mu_{\theta})^{2}}{2\sigma^{2}_{\theta}}-\log(\sqrt{2\pi\sigma^{2}_{\theta}})&\text{if }\textit{lb}_{\textit{pc}}<x<\textit{ub}_{\textit{pc}},\\ \log\left(1-\Phi\left(\frac{\textit{ub}_{\textit{pc}}-\mu_{\theta}}{\sigma_{\theta}}\right)\right)&\text{if }x\geq\textit{ub}_{\textit{pc}}\end{cases}$
B.2 Violation Projection Layer
We define a feasible region of action $x$ as the polyhedron:
$\textit{PH}=\{x\in\mathbb{R}^{n}:Ax\leq b,x\geq 0\},$
where $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$.
Constraints in PH may be violated during optimization. To quantify these violations, we introduce the violation function:
$\mathcal{V}(x)=(Ax-b)_{>0},$
where $\mathcal{V}(x)_{m_{i}}>0$ indicates that constraint ${m_{i}}$ is violated, and $\mathcal{V}(x)_{m_{i}}=0$ means the constraint is satisfied.
Violation Gradient Descent. To minimize constraint violations, we update $x$ for a fixed number of iterations using gradient descent on the violation term $\|\mathcal{V}(x)\|_{2}^{2}$, which represents the squared distance to feasibility. Differentiating with respect to $x$, we derive the update rule:
$x^{\prime}$ $=x-\eta_{v}\nabla_{x}\|\mathcal{V}(x)\|_{2}^{2}$
$=x-\eta_{v}2A^{\top}\mathcal{V}(x).$
Since the step size $\eta_{v}\in(0,1)$ is a tunable parameter, we simplify the update function to:
$x^{\prime}=x-\eta_{v}A^{\top}\mathcal{V}(x).$
Theorem 1 (Convergence of Violation Gradient Descent).
Let $x_{0}\in\mathbb{R}^{n}$ be an initial point, and consider update:
$x_{k+1}=x_{k}-\eta_{v}A^{\top}\mathcal{V}(x_{k}),$
where:
  • $\mathcal{V}(x)=\max(0,Ax-b)$ is the element-wise function onto nonnegative constraint values.
  • $\eta_{v}\in(0,1)$ is a step size parameter.
  • $A\in\mathbb{R}^{m\times n}$ has full row rank.
  • The feasible region $\textit{PH}=\{x\in\mathbb{R}^{n}:Ax\leq b,x\geq 0\}$ is nonempty.
Then, the sequence $\{x_{k}\}$ satisfies:
  1. 1.
    The function $g(x)=\|\mathcal{V}(x)\|_{2}^{2}$ is non-increasing.
  2. 2.
    $x_{k}$ converges to a feasible point $x^{*}$ or a local minimum violation point.
Proof.
Define the violation function:
$g(x)=\|\mathcal{V}(x)\|_{2}^{2}.$
Since $\mathcal{V}(x)$ is an elementwise projection onto nonnegative values, and some function $h(y)=\max(0,y)$ is convex and non-decreasing, then the function $g(x)=\|\mathcal{V}(x)\|_{2}^{2}$ is convex when $Ax-b$ is affine.
We apply gradient descent on $g(x)$ using the update rule:
$x_{k+1}=x_{k}-\eta_{v}\nabla_{x}g(x_{k}).$
By the standard descent lemma Boyd and Vandenberghe [2004], for a sufficiently small step size $\eta_{v}$, we have:
$g(x_{k+1})\leq g(x_{k})-\eta_{v}\|\nabla_{x}g(x_{k})\|_{2}^{2}.$
Since $\eta_{v}>0$ and $\|\nabla_{x}g(x_{k})\|_{2}^{2}\geq 0$, it follows that:
$g(x_{k+1})\leq g(x_{k}).$
Thus, $g(x_{k})$ is non-increasing.
Since $g(x_{k})$ is also lower-bounded by 00, it must converge to some limit $g^{*}\geq 0$. This implies:
$\lim_{k\to\infty}\|\mathcal{V}(x_{k})\|_{2}=\tilde{c},\quad\text{for some }\tilde{c}\geq 0.$
If $\tilde{c}=0$, then $x_{k}$ converges to a feasible point, meaning $\mathcal{V}(x_{k})=0$. If $\tilde{c}>0$, then $x_{k}$ converges to a local minimum of $g(x)$, where no further descent is possible, satisfying $\nabla_{x}g(x_{k})=0,$ which implies $A^{\top}\mathcal{V}(x_{k})=0.$
Appendix C Instance Generator
During training, we simulate problem instances based on a Gaussian distribution with element $i$:
$q^{(i,j,k)}\sim\mathcal{N}\big{(}\mu^{(i,j,k)},\sigma^{(i,j,k)}\big{)}\;\forall(i,j)\in\textit{TR},k\in K.$
Here, $\mu$ is the expected value of cargo demand, initialized by a uniform generator $\mathcal{U}(\textit{lb},\textit{ub})$, while the standard deviation of demand is defined as $\sigma^{(i,j,k)}=\textit{CV}\cdot\mu^{(i,j,k)}$, where the coefficient of variation (CV) controls the spread of each element based on $\mu^{(i,j,k)}$. Note that CV is normally defined as $\textit{CV}^{(i,j,k)}=\frac{\sigma^{(i,j,k)}}{\mu^{(i,j,k)}}$, so we use $\sigma^{(i,j,k)}=\textit{CV}\cdot\mu^{(i,j,k)}$ to control the spread of the distribution.
We initialize $\mu^{(i,j,k)}\sim\mathcal{U}(0,\overline{\mu^{(i,j,k)}})$, where the upper bound on the expected value is found as follows:
$\overline{\mu}=\frac{2\textit{UR}\cdot\mathbf{1}^{\top}c}{\textit{NC}}\qquad,$
where UR is the rate of total utilization present in the demand (e.g., 1.2 means total demand is 120% of total capacity), and $\textit{NC}\in\mathbb{R}^{|TR|}_{>0}$ is a matrix to spread the demand over different elements proportional to the number of transports remaining to be loaded.
During generalization testing, we simulate problem instances based on a continuous uniform generator:
$q^{(i,j,k)}\sim\mathcal{U}(\textit{lb}^{(i,j,k)},\textit{ub}^{(i,j,k)}).$
To ensure similar mean and variance of the instances, we derive parameters $a$ and $b$ from the definition of the continuous uniform distribution, as follows:
$\mu^{(i,j,k)}=(\textit{lb}^{(i,j,k)}+\textit{ub}^{(i,j,k)})/2$
$(\sigma^{(i,j,k)})^{2}=(\textit{ub}^{(i,j,k)}-\textit{lb}^{(i,j,k)})^{2}/12$
We rewrite to:
$\textit{lb}^{(i,j,k)}=\mu^{(i,j,k)}-\sqrt{(12(\sigma^{(i,j,k)})^{2})/2}$
$\textit{ub}^{(i,j,k)}=\mu^{(i,j,k)}+\sqrt{(12(\sigma^{(i,j,k)})^{2})/2}$
Appendix D Multi-stage Stochastic MIP
D.1 Multi-stage Scenario Tree
A scenario tree is a directed tree represented as $T_{\textit{ST}}=(V_{\textit{ST}},E_{\textit{ST}})$, where $V_{\textit{ST}}$ is the set of nodes, each corresponding to a decision or uncertainty realization at a given stage. $E_{\textit{ST}}\subseteq V_{\textit{ST}}\times V_{\textit{ST}}$ is the set of directed edges representing transitions between nodes over time.
The tree consists of:
  1. 1.
    A root node $v_{1}\in V_{\textit{ST}}$, representing the initial state at the first port.
  2. 2.
    Stages $p=1,2,\dots,N_{P}-1$, where each node $v$ belongs to a stage $p(v)$. We denote stages by $p$, as stages are equivalent to ports in a voyage.
  3. 3.
    Branching structure, where each node has child nodes that correspond to possible future realizations.
  4. 4.
    A probability measure $P:V_{\textit{ST}}\to[0,1]$ assigning probabilities to nodes, ensuring:
    $\sum_{v^{\prime}\in\text{child}(v)}\mathbb{P}(v^{\prime})=\mathbb{P}(v),\quad\forall v\in V_{\textit{ST}}.$
  5. 5.
    Scenario paths $\phi\in\mathcal{Z}$, which are root-to-leaf paths representing possible realizations of uncertainty over time.
D.2 MIP Formulation
We define the MPP under demand uncertainty as a multi-stage stochastic MIP.
Decision Variables. The following variables are included:
  • Vessel utilization: $\tilde{u}^{b,d,\phi}_{\textit{tr},k}\in\mathbb{Z}_{\geq 0}$
  • Hatch overstowage: $\tilde{\textit{ho}}^{\phi}_{p,b}\in\mathbb{Z}_{\geq 0}$
  • Makespan of cranes: $\tilde{\textit{cm}}^{\phi}_{p}\in\mathbb{Z}_{\geq 0}$
  • Hatch movement: $\tilde{hm}^{\phi}_{p,b}\in\{0,1\}$
In addition, we introduce the big $\mathbf{M}$ as a sufficiently large constant used to enforce logical constraints.
Objective. The objective function (18) maximizes the revenue and minimizes hatch-overstowage and crane moves costs over scenario paths $\phi\in\mathcal{Z}$ each with probability $\mathbb{P}_{\phi}$. We assume each scenario path has equal probability.
Constraints. Constraint (19) enforces that the onboard utilization cannot exceed the cargo demand $q$, whereas Constraint (20) limits each vessel location to the TEU capacity $c$ for each bay $b\in B$ and deck $d\in D$. In Constraint (21), we indicate that hatches need to be opened if below deck cargo needs to be loaded or discharged. Based on these movements, Constraint (22) models the amount of hatch overstowage in containers. Subsequently, we compute the target of crane moves $\overline{z}$ in Constraint (23), after which Constraint (24) computes the excess number of crane moves $\tilde{\textit{cm}}$.
Additionally, we model the longitudinal and vertical stability in Constraints (26) until (29). First, we compute the longitudinal moment, vertical moment and total weight in Constraints (26), (27) and (25), respectively. Second, Constraint (28) bounds lcg between $\underline{\textit{lcg}}$ and $\overline{\textit{lcg}}$. Third, Constraint (29) bounds vcg between $\underline{\textit{vcg}}$ and $\overline{\textit{vcg}}$. Both lcg and vcg are linearized equivalents of the original Constraints (1) and (2), respectively. Furthermore, we include non-anticipation in Constraint (30) to prevent leveraging future demand realizations.
max $\sum_{\phi\in\mathcal{Z}}\mathbb{P}_{\phi}\sum_{p\in{P}}\sum_{b\in{B}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{+}}(p)}f_{1}\tilde{u}^{b,d,\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\qquad\qquad-f_{2}\tilde{\textit{ho}}^{\phi}_{p,b}-f_{3}\tilde{\textit{cm}}^{\phi}_{p}$ (18)
s.t. $\sum_{b\in{B}}\sum_{d\in{D}}\tilde{u}^{b,d,\phi}_{\textit{tr},k}\leq q^{\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\forall p\in{P},\;\textit{tr}\in\textit{TR}^{\textit{OB}}(p),\;k\in{K},\;\phi\in\mathcal{Z}$ (19)
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{OB}}(p)}\textit{teu}_{\textit{tr},k}\tilde{u}^{b,d,\phi}_{\textit{tr},k}\leq c_{b,d}$
$\qquad\qquad\qquad\forall p\in{P},\;b\in{B},\;d\in{D},\;\phi\in\mathcal{Z}$ (20)
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}\tilde{u}^{b,d_{h},\phi}_{\textit{tr},k}\leq\mathbf{M}\tilde{hm}^{\phi}_{p,b}$
$\qquad\qquad\qquad\forall p\in{P},\;b\in{B},\;\phi\in\mathcal{Z}$ (21)
$\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{\textit{AC}}(p)}\tilde{u}^{b,d_{o},\phi}_{\textit{tr},k}-M(1-\tilde{hm}^{\phi}_{p,b})\leq\tilde{\textit{ho}}^{\phi}_{p,b}$
$\qquad\qquad\qquad\forall p\in{P},\;b\in{B},\;\phi\in\mathcal{Z}$ (22)
${\overline{z}}^{\phi}_{p}=(1+\delta^{\textit{cm}})\frac{2}{|{B}|}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}\sum_{k\in{K}}q^{\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\forall p\in{P},\;\phi\in\mathcal{Z}$ (23)
${\overline{z}}^{\phi}_{p}-\sum_{b\in{b^{\prime}}}\sum_{d\in{D}}\sum_{k\in{K}}\sum_{\textit{tr}\in\textit{TR}^{M}_{p}}\tilde{u}^{b,d,\phi}_{\textit{tr},k}\leq\tilde{\textit{cm}}^{\phi}_{p}$
$\qquad\qquad\qquad\forall p\in{P},\;{b^{\prime}}\in{B}^{\prime},\;\phi\in\mathcal{Z}$ (24)
$\textit{tw}^{\phi}_{p}=\sum_{k\in{K}}w_{k}\sum_{\textit{tr}\in\textit{TR}^{\textit{OB}}(p)}\sum_{d\in{D}}\sum_{b\in{B}}\tilde{u}^{b,d,\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\forall p\in{P},\;\phi\in\mathcal{Z}$ (25)
$\textit{lm}^{\phi}_{p}=\sum_{b\in{B}}\textit{ld}_{b}\sum_{k\in{K}}w_{k}\sum_{\textit{tr}\in\textit{TR}^{\textit{OB}}(p)}\sum_{d\in{D}}\tilde{u}^{b,d,\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\forall p\in{P},\;\phi\in\mathcal{Z}$ (26)
$\textit{vm}^{\phi}_{p}=\sum_{d\in{D}}\textit{vd}_{d}\sum_{k\in{K}}w_{k}\sum_{\textit{tr}\in\textit{TR}^{\textit{OB}}(p)}\sum_{b\in{B}}\tilde{u}^{b,d,\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\forall p\in{P},\;\phi\in\mathcal{Z}$ (27)
$\underline{\textit{lcg}}\textit{tw}^{\phi}_{p}\leq\textit{lm}^{\phi}_{p}\leq\overline{\textit{lcg}}\textit{tw}^{\phi}_{p}$
$\qquad\qquad\qquad\forall p\in{P},\;\phi\in\mathcal{Z}$ (28)
$\underline{\textit{vcg}}\textit{tw}^{\phi}_{p}\leq\textit{vm}^{\phi}_{p}\leq\overline{\textit{vcg}}\textit{tw}^{\phi}_{p}$
$\qquad\qquad\qquad\forall p\in{P},\;\phi\in\mathcal{Z}$ (29)
$\tilde{u}^{b,d,\phi^{\prime}}_{\textit{tr},k}=\tilde{u}^{b,d,\phi}_{\textit{tr},k}$
$\qquad\qquad\qquad\forall p\in{P},\textit{tr}\in\textit{TR}^{\textit{L}}(p),k\in{K},$
$\qquad\qquad\qquad b\in{B},d\in{D},\phi,\phi^{\prime}\in\mathcal{Z}\mid q^{\phi}_{[p-1]}=q^{\phi^{\prime}}_{[p-1]}$ (30)
Appendix E Deep RL Implementation Details
To prevent confusion, we
E.1 PPO Algorithm
PPO is an on-policy reinforcement learning algorithm that seeks to maximize expected cumulative reward while enforcing stable policy updates via clipped importance sampling Schulman et al. [2017], as outlined in Algorithm 3. The agent collects trajectories, computing $n_{\text{ppo}}$-step return to evaluate performance with $V_{\theta}(s)$ as estimated state value:
$G_{t}^{(n_{\text{ppo}})}=\sum_{k_{\text{ppo}}=0}^{n_{\text{ppo}}-1}\gamma^{k}r_{t+k_{\text{ppo}}}+\gamma^{n_{\text{ppo}}}V_{\theta}(s_{t+n_{\text{ppo}}}),$ (31)
To reduce variance, we adopt Generalized Advantage Estimation (GAE) Schulman et al. [2016]:
$\hat{A}_{t}^{\text{GAE}}$ $=\sum_{l_{\text{ppo}}=0}^{\infty}(\gamma\lambda)^{l_{\text{ppo}}}\delta_{t+l_{\text{ppo}}},$ (32)
$\delta_{t}$ $=r_{t}+\gamma V_{\theta}(s_{t+1})-V_{\theta}(s_{t}).$ (33)
Here, $\delta_{t}$ is the temporal difference (TD) residual, which quantifies the advantage of taking action $x_{t}$ at state $s_{t}$.
The actor is updated using the PPO clipped surrogate loss:
$\mathcal{L}_{\text{actor}}(\theta)$ $=\mathbb{E}_{t}\Big{[}\min\big{(}\text{ratio}_{t}(\theta)\hat{A}_{t}^{\text{GAE}},$
$\quad\text{clip}(\text{ratio}_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}^{\text{GAE}}\big{)}\Big{]},$ (34)
where the probability ratio is defined as:
$\text{ratio}_{t}(\theta)=\frac{\pi_{\theta}(x_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(x_{t}|s_{t})}.$ (35)
The critic aims to minimize the squared TD error:
$\mathcal{L}_{\text{critic}}(\theta)=\mathbb{E}_{t}\Big{[}\big{(}V_{\theta}(s_{t})-G_{t}^{(n_{\text{ppo}})}\big{)}^{2}\Big{]}.$ (36)
Finally, the total PPO objective, including feasibility regularization from Equation (11), is given by:
$\mathcal{L}(\theta)$ $=\mathcal{L}_{\text{actor}}(\theta)+\lambda_{c}\mathcal{L}_{\text{critic}}(\theta)+\lambda_{f}\mathcal{L}_{\text{feas}}(\theta)$
$\qquad-\lambda_{e}\mathbb{E}_{t}\big{[}\text{entropy}(\pi_{\theta})\big{]},$ (37)
where $\lambda_{c}$, $\lambda_{f}$, and $\lambda_{e}$ are weighting coefficients for the critic loss, feasibility regularization, and entropy regularization respectively.
Algorithm 3 Proximal Policy Optimization (PPO)
0:  Model parameters $\theta$, steps $n$, learning rate $\eta$
1:  for each gradient update do
2:     for each step $t$ do
3:        Collect $n$-step trajectories $\{(s_{t},x_{t},r_{t},s_{t+1})\}$
4:        Compute $n$-step returns $G_{t}^{(n_{\text{ppo}})}$
5:        Compute advantage estimates $\hat{A}_{t}^{\text{GAE}}$
6:     end for
7:     Update parameters: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{L}(\theta)$
8:  end for
9:  return  Policy $\pi_{\theta}$
E.2 SAC Algorithm
Soft Actor-Critic (SAC) is an off-policy reinforcement learning algorithm that optimizes both reward maximization and entropy to encourage efficient exploration Haarnoja et al. [2018], as outlined in Algorithm 4. It is based on maximum entropy reinforcement learning, which aims to learn a stochastic policy that not only maximizes cumulative rewards but also maintains high entropy for robustness and stability. SAC leverages a soft Q-learning approach, using two Q-functions to mitigate overestimation bias, an entropy-regularized policy update, and an automatically adjusted temperature parameter to balance exploration and exploitation.
The algorithm maintains an actor network for policy learning, two Q-function critics for value estimation, a target Q-network for stable learning, and an adaptive temperature parameter to regulate entropy. The loss functions for standard SAC are derived from the Bellman backup equation and the policy gradient formulation, ensuring convergence to an optimal stochastic policy. We also include feasibility regularization from Equation (11) in the actor loss.
  • Compute target Q-value:
$Q_{\text{target}}(s_{t},x_{t})$ $=r_{t}+\gamma\mathbb{E}_{s_{t+1},x_{t+1}\sim\pi}\Big{[}$
$\min_{l=1,2}Q_{\theta}^{l}(s_{t+1},x_{t+1})-\alpha\log\pi_{\theta}(x_{t+1}|s_{t+1})\Big{]}$
  • Critic loss:
    $\mathcal{L}_{\text{critic}}(\theta)=\mathbb{E}\Big{[}(Q_{\theta}(s_{t},x_{t})-Q_{\text{target}}(s_{t},x_{t}))^{2}\Big{]}$
  • Actor loss:
    $\mathcal{L}_{\text{actor}}(\theta)=\mathbb{E}\Big{[}\alpha\log\pi_{\theta}(x_{t}|s_{t})-Q_{\theta}(s_{t},x_{t})+\lambda_{f}\mathcal{L}_{\text{feas}}(\theta)\Big{]}$
  • Temperature loss:
    $\mathcal{L}_{\alpha}(\theta)=\mathbb{E}\Big{[}-\alpha(\log\pi_{\theta}(x_{t}|s_{t})+\text{entropy}_{\text{target}})\Big{]}$
This formulation ensures stability and encourages exploration by adapting the trade-off between exploitation and exploration dynamically.
Algorithm 4 Soft Actor-Critic (SAC)
0:  Parameters: actor $\theta_{\text{actor}}$, critics $\theta_{\text{critic}}^{1},\theta_{\text{critic}}^{2}$, targets $(\theta_{\text{target}}^{1},\theta_{\text{target}}^{2})=(\theta_{\text{critic}}^{1},\theta_{\text{critic}}^{2})$, temperature $\alpha$, learning rate actor $\eta_{a}$, learning rate critic $\eta_{c}$, learning rate temperature $\eta_{\alpha}$, soft update parameter $\tau$, replay buffer $\mathcal{D}$.
1:  for each step $t$ do
2:     Sample action $x_{t}\sim\pi_{\theta}(x_{t}|s_{t})$
3:     Perform transition $s_{t+1}\sim\mathcal{T}(s_{t+1}|s_{t},x_{t})$
4:     Observe reward $r_{t}=\mathcal{R}(s_{t},x_{t})$,
5:     Store $(s_{t},x_{t},r_{t},s_{t+1})$ in $\mathcal{D}$.
6:  end for
7:  for each gradient update do
8:     Sample a minibatch $(s_{t},x_{t},r_{t},s_{t+1})$ from $\mathcal{D}$.
9:     Compute target Q-value: $Q_{\text{target}}(s_{t},x_{t})$
10:     Update parameters:  $\theta_{\text{critic}}^{l}\leftarrow\theta_{\text{critic}}^{l}-\eta_{c}\nabla_{l}\mathcal{L}_{\text{critic}}(\theta)\text{ for }l\in\{1,2\}$$\theta_{\text{actor}}\leftarrow\theta_{\text{actor}}-\eta_{a}\nabla\mathcal{L}_{\text{actor}}(\theta)$$\alpha\leftarrow\alpha-\eta_{\alpha}\nabla\mathcal{L}_{\alpha}(\theta)$$\theta_{\text{target}}^{l}\leftarrow\tau\theta_{\text{critic}}^{l}+(1-\tau)\theta_{\text{target}}^{l}\text{ for }l\in\{1,2\}$
11:  end for
E.3 Hyperparameters
Parameters of the MPP environment are shown in Table 3.
Parameters Symbol Value
Voyage length $N_{P}$ 4
Number of bays $N_{B}$ 10
Cardinality deck set $|D|$ 2
Cardinality cargo set $|K|$ 12
[…]
Vessel TEU $\mathbf{1}^{\top}c$ 1,000
Long term contract reduction LR 0.3
Utilization rate demand UR 1.1
lcg bounds $(\underline{\textit{lcg}},\overline{\textit{lcg}})$ (0.85,1.05)
vcg bounds $(\underline{\textit{vcg}},\overline{\textit{vcg}})$ (0.95,1.15)
Crane moves allowance $\delta^{\textit{cm}}$ 0.25
Overstowage costs $\textit{ct}^{\textit{ho}}$ 0.33
Crane move costs $\textit{ct}^{\textit{ho}}$ 0.5
Table 3: Environment parameters.
Table 4 provides the hyperparameters of projected and vanilla PPO and SAC.
Settings Projection Algorithms Vanilla Algorithms
Hyperparameters Symbol PPO SAC PPO SAC
Actor Network Attention Attention Attention Attention
Number of Heads 8 8 4 4
Hidden Layer Size 128 128 256 256
Encoder Layers 3 3 2 1
Decoder Layers 3 3 3 3
Critic Network $1\times\text{MLP}$ $2\times\text{MLP}$ $1\times\text{MLP}$ $2\times\text{MLP}$
Critic Layers 4 4 4 4
Target Network No Soft Update No Soft Update
Target Update Rate $\tau$ N/A 0.005 N/A 0.005
Dropout Rate $0.009$ $0.009$ $0.073$ $0.164$
Max Policy Std. $1.931$ $1.931$ $1.118$ $1.779$
Optimizer Adam Adam Adam Adam
Learning Rate $\eta$ $2.04\times 10^{-4}$ $2.04\times 10^{-4}$ $9.64\times 10^{-4}$ $9.10\times 10^{-4}$
Batch Size 64 64 64 64
Embedding Size 128 128 128 128
Discount Factor $\gamma$ 0.99 0.99 0.99 0.99
GAE $\lambda$ 0.95 N/A 0.95 N/A
Value Coefficient $\lambda_{c}$ $0.50$ N/A $0.50$ N/A
Entropy Coefficient $\lambda_{e}$ $0.010$ Learned $0.061$ Learned
Feasibility Penalty $\lambda_{f}$ $0.0677$ $0.0677$ $0.302$ $0.065$
Clip Parameter $\epsilon$ 0.2 N/A 0.2 N/A
Replay Buffer No Yes No Yes
Replay Buffer Size N/A $10^{4}$ N/A $10^{4}$
Mini-batch Size 16 16 32 16
Update Epochs 5 1 1 1
Entropy Target N/A $-|{X}|$ N/A $-|{X}|$
Projection Learning Rate $\eta_{v}$ $0.05$ $0.05$ N/A N/A
Projection Epochs $100$ $100$ N/A N/A
Inference Projection Stop $\delta_{v}$ $100$ $100$ N/A N/A
Training Budget $7.2\times 10^{7}$ $7.2\times 10^{7}$ $7.2\times 10^{7}$ $7.2\times 10^{7}$
Validation Budget $5.0\times 10^{3}$ $5.0\times 10^{3}$ $5.0\times 10^{3}$ $5.0\times 10^{3}$
Validation Frequency Every 20% Every 20% Every 20% Every 20%
Table 4: Comparison of hyperparameters for projected and vanilla PPO and SAC.
E.4 Additional Experiments
In Table 5, we analyze different configurations of feasibility regularization (FR). First, we evaluate performance under the same hyperparameters as AM-P policies. Second, we examine the effect of significantly increasing $\lambda_{f}$. Third, we assess performance with specific hyperparameter tuning, including $\lambda_{f}$. These experiments indicate that FR can reduce the distance to the feasible region, however, achieving fully feasible instances remains a challenge.
Methods Testing ($\boldsymbol{N=30}$)
Alg. Model F.M. H.P. $\lambda_{f}$ Ob. (\$) Time (s) F.I. (%) $d(\textit{PH}(s_{t}))$
SAC AM FR Proj. 0.0677 1139.78 13.82 0.00 62.77
SAC AM FR Proj. 0.677 1031.55 13.45 0.00 120.74
SAC AM FR Tune 0.065 1113.03 12.63 0.00 37.55
PPO AM FR Proj. 0.0677 2606.87 13.36 0.00 3171.86
PPO AM FR Proj. 0.677 2593.29 13.20 0.00 3381.59
PPO AM FR Tune 0.302 1842.46 11.74 0.00 754.21
Table 5: Performance evaluation on $N$ instances of feasibility regularization (FR) with hyperparameter (H.P.) settings: projected hyperparameters (Proj.), ensuring a fair comparison with projection-based policies, and tuned hyperparameter (Tune), optimized specifically for PPO and SAC with FR. While we use $\lambda_{f}$ as the control parameter for FR, tuning involves adjusting a Lagrangian multiplier for each constraint. Average performance metrics include objective value in profit (Ob.), inference time in seconds (Time), percentage of feasible instances (F.I.), and total absolute distance to the feasible region $d(\textit{PH}(s_{t}))$. Note that † indicates infeasible objectives.